Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Leila is playing a carnival game in which she is given 4

Expert replies
by AAPL » Tue Jan 29, 2019 3:48 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

Manhattan Prep

Leila is playing a carnival game in which she is given 4 chances to throw a ball through a hoop. If her of success on each throw is 1/5, what is the chance that she will succeed on at least 3 of the throws?

A. 1/5^4
B. 1/5^3
C. 6/5^4
D. 13/5^4
E. 17/5^4

OA E
Join the discussion
Source: — Problem Solving |

by fskilnik@GMATH » Tue Jan 29, 2019 4:00 am
AAPL wrote:Manhattan Prep

Leila is playing a carnival game in which she is given 4 chances to throw a ball through a hoop. If her of success on each throw is 1/5, what is the chance that she will succeed on at least 3 of the throws?

A. 1/5^4
B. 1/5^3
C. 6/5^4
D. 13/5^4
E. 17/5^4
$$P\left( {{\rm{success}}\,\,{\rm{in}}\,\,{\rm{any}}\,\,{\rm{throw}}} \right) = {1 \over 5}$$
$$? = P\left( {{\rm{success}}\,\,{\rm{in}}\,{\rm{all}}\,\,4\,\,{\rm{throws}}} \right) + P\left( {{\rm{success}}\,\,{\rm{in}}\,{\rm{exactly}}\,\,3\,\,{\rm{throws}}} \right)$$
$$\left. \matrix{
P\left( {{\rm{success}}\,\,{\rm{in}}\,{\rm{all}}\,\,4\,\,{\rm{throws}}} \right)\,\,\,\mathop = \limits^{{\rm{independency}}} \,\,\,{\left( {{1 \over 5}} \right)^4} = {1 \over {{5^4}}} \hfill \cr
P\left( {{\rm{success}}\,\,{\rm{in}}\,{\rm{exactly}}\,\,3\,\,{\rm{throws}}} \right)\,\,\,\mathop = \limits^{{\rm{independency}}} \,\,\,C\left( {4,3} \right) \cdot {\left( {{1 \over 5}} \right)^3} \cdot {4 \over 5}\,\,\, = \,\,\,{{4 \cdot 4} \over {{5^4}}}\,\,\,\, \hfill \cr} \right\}\,\,\,\,\, \Rightarrow \,\,\,\,\,? = {{1 + 16} \over {{5^4}}}\,\,\,\,\,\,\, \Rightarrow \,\,\,\,\,\,\,\left( {\rm{E}} \right)$$

We follow the notations and rationale taught in the GMATH method.

Regards,
Fabio.
Fabio Skilnik :: GMATH method creator ( Math for the GMAT)
English-speakers :: https://www.gmath.net
Portuguese-speakers :: https://www.gmath.com.br
Join the discussion

by Brent@GMATPrepNow » Tue Jan 29, 2019 5:44 am
AAPL wrote:Leila is playing a carnival game in which she is given 4 chances to throw a ball through a hoop. If her of success on each throw is 1/5, what is the chance that she will succeed on at least 3 of the throws?

A. 1/5^4
B. 1/5^3
C. 6/5^4
D. 13/5^4
E. 17/5^4
OA E
Given: P(succeeds on 1 throw) = 1/5

P(succeeds at least 3 times) = P(succeeds 4 times OR succeeds 3 times)
= P(succeeds 4 times) + P(succeeds 3 times)

P(succeeds 4 times)
P(succeeds 4 times) = P(succeeds 1st time AND succeeds 2nd time AND succeeds 3rd time AND succeeds 4th time)
= P(succeeds 1st time) x P(succeeds 2nd time) x P(succeeds 3rd time) x P(succeeds 4th time)
= 1/5 x 1/5 x 1/5 x 1/5
= 1/5�

P(succeeds 3 times)
Let's examine one possible scenario in which Leila succeeds exactly 3 times:
P(FAILS the 1st time AND succeeds 2nd time AND succeeds 3rd time AND succeeds 4th time)
= P(FAILS the 1st time) x P(succeeds 2nd time) x P(succeeds 3rd time) x P(succeeds 4th time)
= 4/5 x 1/5 x 1/5 x 1/5
= 4/5�
Keep in mind that this is only ONE possible scenario in which Leila succeeds exactly 3 times (Leila fails the 1st time).
Leila can also FAIL the 2nd time, or the 3rd time or the 4th time.
Each of these probabilities will also equal 4/5³
So, P(succeeds 3 times) = 4/5� + 4/5� + 4/5� + 4/5�
= 16/5�

So, P(succeeds AT LEAST 3 times) = P(succeeds 4 times) + P(succeeds 3 times)
= 1/5� + 16/5�
= 17/5�

Answer: E


Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by swerve » Tue Jan 29, 2019 11:32 am
P (that she succeeds) = 1/5
P(of success exactly 3 times) = 1/5*1/5*1/5*4/5 = 4/5^4. Now, these can occur in 4!/3!*1! = 4. hence, P(success 3 times)= 4*4/5^4
P (success all 4 times)= 1/5^4
Answer= 16/5^4+1/5^4 = 17/5^4.

Regards!
Join the discussion

by Scott@TargetTestPrep » Thu Jan 31, 2019 5:36 pm
AAPL wrote:Manhattan Prep

Leila is playing a carnival game in which she is given 4 chances to throw a ball through a hoop. If her of success on each throw is 1/5, what is the chance that she will succeed on at least 3 of the throws?

A. 1/5^4
B. 1/5^3
C. 6/5^4
D. 13/5^4
E. 17/5^4

OA E
We need to determine the probability that Leila succeeds on exactly 3 throws OR on all 4 throws.

Scenario 1: succeeds on exactly 3 throws

We can let Y denote a successful throw and N denote a non-successful throw:

P(Y-Y-Y-N) = 1/5 x 1/5 x 1/5 x 4/5 = 4/(5^4)

However, we must account for the order of Y-Y-Y-N. Using our formula for indistinguishable items, Y, Y, Y, and N can be arranged in 4!/3! = 4 ways.

Thus, the probability of succeeding on exactly 3 throws (out of 4 attempts) is 4/(5^4) x 4 = 16/(5^4).

Scenario 2: succeeds on all 4 attempts

P(Y-Y-Y-Y) = 1/5 x 1/5 x 1/5 x 1/5 = 1/(5^4)

Thus, the probability of succeeding on 3 throws out of 4 or 4 throws out of 4 is 16/(5^4) + 1/(5^4) = 17/5^4.

Answer: E

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion